OpenStudy (isabel☺):

Solve the DE (x+2y-1)dx- (2x+y-5)dy=0

OpenStudy (isabel☺):

by inspection?

ganeshie8 (ganeshie8):

did u check if it is exact ?

OpenStudy (isabel☺):

uhmm..it is exact.. but we're (told) supposed to solve it by either inspection, by obvious subs., by formulas, others .. not by the previous lesson like exact :)

OpenStudy (isabel☺):

please help me @ganeshie8

ganeshie8 (ganeshie8):

can you check if the sign sign between differentials is really "-" ?

ganeshie8 (ganeshie8):

(x+2y-1)dx\(\color{red}{-}\) (2x+y-5)dy=0

ganeshie8 (ganeshie8):

i think that should be "+", can you please double check

OpenStudy (isabel☺):

wait

ganeshie8 (ganeshie8):

it wont be exact if that sign is "-"

OpenStudy (isabel☺):

it's plus..

OpenStudy (isabel☺):

i'm gonna try to solve it too

OpenStudy (isabel☺):

nope...

ganeshie8 (ganeshie8):

nope for what ?

OpenStudy (isabel☺):

it is a minus sign

ganeshie8 (ganeshie8):

Oh its okay, so you want to work it by substitution is it ?

ganeshie8 (ganeshie8):

here is the same question, let me know if soemthing doesn't make sense in this : http://math.stackexchange.com/questions/448839/how-to-solve-x-2y-4dx-2x-y-5dy-0

OpenStudy (isabel☺):

coefficient of linear : intersecting lines

ganeshie8 (ganeshie8):

yes try substituting : \[x = X-a \\ y = Y-b\]

ganeshie8 (ganeshie8):

choose \(a\) and \(b\) valuse such that they eliminate the constants in the original equaiton

OpenStudy (isabel☺):

we use : x=u+h y=v+k

ganeshie8 (ganeshie8):

anything is fine :)

OpenStudy (isabel☺):

(h,k)=(3,-1)

ganeshie8 (ganeshie8):

perfect !

OpenStudy (isabel☺):

so x=y+3 y=v-1 ??

OpenStudy (isabel☺):

i mean x=u+ 3

ganeshie8 (ganeshie8):

yes, x = u + 3 y = v - 1

ganeshie8 (ganeshie8):

resultant equation would be : \[\large (u+2v)du - (2u+v)dv = 0\] ?

OpenStudy (isabel☺):

yes, mine too :)

OpenStudy (isabel☺):

and then?

ganeshie8 (ganeshie8):

let \(\large v = tu\) \(\large dv = tdu + udt\)

OpenStudy (isabel☺):

homogeneous?

ganeshie8 (ganeshie8):

\[\large (u+2v)du - (2u+v)dv = 0\] becomes \[\large (u+tu)du - (2u+tu)(tdu + udt ) = 0\] \[\large (1+t)du - (2+t)(tdu + udt ) = 0\] \[\large (1+t-t(1+t))du - u(2+t) dt = 0\] \[\large (1-t^2)du - u(2+t) dt = 0\] \[\large (1-t^2)du = u(2+t) dt \] \[\large \dfrac{du}{u} = \dfrac{2+t}{1-t^2} dt \]

ganeshie8 (ganeshie8):

verify if the calculaitons are correct or not :) if everything is right, you may integrate both sides

OpenStudy (isabel☺):

\[C= \ln(x-3) + 3/2 \ln(1-\frac{ y+1 }{ x-3 }) - 1/2 \ln(1+ \frac{ y+1 }{ x-3 })\] ???

ganeshie8 (ganeshie8):

something like that... just check the signs again ^^

OpenStudy (isabel☺):

THANKS!!!!